Tyson is 9 times the age of his daughter, Charlotte. In 5 years, he will be 4 times the age of Charlotte. How old are they now?
Answer by Theo(13342)


T equals tyson's age.
C = charlotte's age.
T = 9C
T + 5 = 4 * (C + 5)

simplify second equation to get:
T + 5 = 4C + 20
replace T with 9C to get:
9C + 5 = 4C + 20
subtract 4C from both sides of the equation and subtract 5 from both sides of the equation to get:
5C = 15
solve for C to get C = 3
you have C = 3 and T = 9 * 3 = 27
in 5 years, C will be equal to 8 and T will be equal to 32.
T will therefore be equal to 4 * C.
this confirms the ages of Tyson and Charlotte are correct.
tyson is 27 and charlotte is 3.
tyson's age is 9 * charlotte's age.
in 5 years, tyson will be 32 and charlotte will be 8.
tyson's age will be 4 * charlotte's age.